### v1

# Pulse-train propagation in nonlinear Kerr media governed by higher-order dispersion

Abstract
We discover three novel classes of pulse-train waveforms in an optical Kerr nonlinear medium possessing all orders of dispersion up to the fourth order. We show that both single- and double-humped pulse-trains can be formed in the nonlinear medium. A distinguishing property is that these structures have different amplitudes, widths and wavenumbers but equal velocity which depends on the three dispersion parameters. More importantly, we find that the relation between the amplitude and duration of all the newly obtained pulse-trains is determined by the sign of a joint parameter solely. The results show that those optical waves are general, in the sense that no specified conditions on the material parameters are assumed. Considering the long-wave limit, the derived pulse-trains degenerate to soliton pulses of the quartic and dipole kinds.
pacs:
05.45.Yv, 42.65.Tg
I Introduction
The transmission of soliton pulses in nonlinear media has been the focus of continuing interest, stimulated by their important potential applications in long-distance optical communication systems and optical switching devises [1]. The formation of such wave packets in a monomode optical fiber results from the exact balance between the pulse-spreading effects of group-velocity dispersion and self-phase modulation effects due to the intensity-dependent refractive index of the fiber [2]. Experimental observation of solitons have been reported in various optical systems, including single-mode optical fibers [3], femtosecond lasers [4] and bulk optical materials [5].
In the picosecond regime, the propagation of light pulses inside a single-mode optical fiber was shown to be governed by the nonlinear Schrödinger equation (NLSE), which is commonly derived with use of the slowly varying envelope approximation [8]. This physically relevant model allows for either bright or dark solitons in the anomalous and normal dispersion regimes, respectively [9]. However, for much shorter light pulses, there is also appearance of additional higher-order effects in the optical system. For example, the third-order dispersion significantly influence the short pulses whose widths are nearly 50 femtosecond as they propagate through the nonlinear medium [10, 11]. The fourth-order dispersion is also found to be important when pulses are shorter than 10 femtosecond [10, 11]. In such a situation, extensions of the NLSE model that take into account the influence of diverse physical phenomena on short-pulse propagation and generation are therefore essential to describe the wave dynamics accurately. One notes that higher-order effects introduce several new important phenomena into the system dynamics that are absent in Kerr nonlinear medium.
A great interest has been given recently to the generation of pulse-trains in various physical systems, including birefringent optical fibers [12, 13, 14, 15, 16], negative-index materials [17], Bose-Einstein condensates [18], etc. It can be relevant to note that such kind of pulse shape is interesting in itself as another type of physically relevant nonlinear waves in optical fibers besides the soliton-type waveform. These structures exhibit many interesting properties that make them potentially in problems of signal propagation in nonlinear fibers [19]. It is intersting to note that pulse-trains can be produced experimentally. For example, the formation of cnoidal wave has been recently demonstrated in a fiber laser either in the net normal or net anomalous cavity dispersion regime [20]. Importantly, such arrays of pulses may be created by the modulational instability of the continuous-wave signal [21], or from a dual-frequency pump [22]. In optics, such nonlinear waves play a pivotal role in the analysis of the data transmission in fiber-optic telecommunications links [23].
Identification of new kinds of pulse-trains in optical systems is an issue worth considering. It is pertinent to mention that the finding of such structures may lead to novel dynamical behaviors of waves and opens the possibility for experimental investigations of various wave phenomena. In this work, we demonstrate for the first time that an optical Kerr nonlinear medium exhibiting all orders of dispersion up to the fourth order supports the existence of three novel classes of pulse-trains which take either single- or double-hump forms.
The paper is structured as follows. Section II provides the analytic framework necessary to find the traveling wave solutions of the extended NLSE model that governs the transmission of extremely short pulses throught an optical fiber medium exhibiting all orders of dispersion up to the fourth order. In Section III, we derive analytical pulse-train solutions of the model and their characteristics. We also discuss here the propagation properties of obtained pulse-trains in the fiber medium. Finally, some conclusions are given in Sec. VI.
II Model and traveling waves
We consider the evolution of an optical pulse in a highly dispersive fiber medium having a Kerr nonlinearity. The governing model for the slowly varying envelope $\psi$ of the optical pulse can be described by the extended NLSE [6, 7],
$i\frac{\partial\psi}{\partial z}=\alpha\frac{\partial\psi^{2}}{\partial\tau^{2}}+i\sigma\frac{\partial\psi^{3}}{\partial\tau^{3}}-\epsilon\frac{\partial\psi^{4}}{\partial\tau^{4}}-\gamma\left|\psi\right|^{2}\psi.$ (1)
where $\tau=t-\beta_{1}z$ is the retarded time, $z$ is the longitudinal coordinate, $\alpha=\beta_{2}/2$, $\sigma=\beta_{3}/6$, $\epsilon=\beta_{4}/24$, and $\gamma$ is the nonlinear parameter. Here the parameters $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ are the first-, second-, third-, and fourth-order dispersion coefficients, respectively.
Theoretical investigations based on Eq. (1) are mainly concentrated on the existence and propagation properties of localized pulses due to their fundamental importance in the understanding of various physical phenomena in the system. For example, Kruglov and Harvey [6] demonstrated the existence of stable solitary waves with a $\mathrm{sech}^{2}$ shape for the model. Triki and Kruglov [7] discussed the dipole soliton solution of Eq. (1) in the presence the inhomogeneities of media. Cavalcanti et al. [24] analyzed the modulational instability phenomena of Eq. (1) in the region of the minimum group-velocity dispersion. Karpman et al. [25, 26] investigated the time behavior of the amplitudes, velocities, and other parameters of radiating solitons. Shagalov [27] studied the effect of the third- and fourth-order dispersions on the modulational instability. Roy et al. [29] discussed the role of third- and fourth-order dispersions in the radiation emitted by fundamental soliton pulses in the form of dispersive waves within the framework of the dimensionless form of the model (1). Kruglov [28] obtained the periodic wave solutions that take the $\mathrm{cn}^{2}$ shape for this governing equation [28]. In what follows, we present three novel exact pulse-train solutions with interesting properties for the extended NLSE (1), which are obtained without necessarily assuming a specified condition on the fiber parameters.
In order to find some interesting analytical solutions of the model (1), we consider a solution of the form
$\psi(z,\tau)=\rho(\xi)\exp[i(\kappa z-\delta\tau+\theta)],$ (2)
where $\rho(\xi)$ is a real function, $\xi=\tau-qz$ is a traveling coordinate with $q=v^{-1}$ being the inverse velocity. Also, $\kappa$ and $\delta$ are the respective real parameters representing the wavenumber and frequency shift, while $\theta$ represents the phase of the pulse at $z=0$.
Substitution of Eq. (1) into the extended NLSE (2) leads to a set of coupled differential equations:
$(\sigma+4\epsilon\delta)\frac{d^{3}\rho}{d\xi^{3}}+(q-2\alpha\delta-3\sigma\delta^{2}-4\epsilon\delta^{3})\frac{d\rho}{d\xi}=0,$ (3)
$\epsilon\frac{d^{4}\rho}{d\xi^{4}}-(\alpha+3\sigma\delta+6\epsilon\delta^{2})\frac{d^{2}\rho}{d\xi^{2}}+\gamma\rho^{3}-(\kappa-\alpha\delta^{2}-\sigma\delta^{3}-\epsilon\delta^{4})\rho=0.$ (4)
Here, we refer the solution of the extended NLSE where the function $\psi(z,\tau)$ is given by Eq. (2) with $\rho(\xi)\neq\mathrm{constant}$ as non-plain wave or traveling wave solution.
The system of Eqs. (3) and (4) with $\epsilon\neq 0$ yields the non-plain wave solutions if and only if the next relations are satisfied:
$q=2\alpha\delta+3\sigma\delta^{2}+4\epsilon\delta^{3},~{}~{}~{}~{}\delta=-\frac{\sigma}{4\epsilon}.$ (5)
Note that Eq. (3) is satisfied for an arbitrary function $\rho(\xi)$ according to the conditions in Eq. (5) with $\epsilon\neq 0$. The relations in Eq. (5) leads to the next expression for the velocity $v=1/q$ defined in the retarded frame,
$v=\frac{8\epsilon^{2}}{\sigma(\sigma^{2}-4\alpha\epsilon)}.$ (6)
The latter relation shows that the dispersion parameters $\beta_{i}$ $(i=2,3,4)$ included in the parameters $\alpha$, $\sigma$ and $\epsilon$ have a major influence on the wave velocity. This implies that the velocity of the pulse can be controlled by an appropriate manipulation of the dispersion parameters.
With use of the relations in Eq. (5), the system of Eqs. (3) and (4) reduces to the ordinary nonlinear differential equation,
$\epsilon\frac{d^{4}\rho}{d\xi^{4}}+b\frac{d^{2}\rho}{d\xi^{2}}-c\rho+\gamma\rho^{3}=0,~{}~{}~{}~{}~{}~{}~{}$ (7)
where the parameters $b$ and $c$ are
$b=\frac{3\sigma^{2}}{8\epsilon}-\alpha,~{}~{}~{}~{}c=\kappa+\frac{\sigma^{2}}{16\epsilon^{2}}\left(\frac{3\sigma^{2}}{16\epsilon}-\alpha\right).$ (8)
Equation (7) describes the evolution of the field amplitude in the highly dispersive optical fiber. Now the natural question is whether the amplitude equation (7) possesses exact pulse-train solutions without any specified condition on the coefficients or not. In following, we have identified novel families of pulse-train solutions with some interesting properties for this evolutional equation in presence of all physical parameters.
III Results and discussion
In this section, we demonstrate that new kinds of pulse-trains can exist in the fiber system governed by Eq. (1). Remarkably, we find that these propagating waves possess distinct amplitudes and consequently different powers, which is highly desired from a practical point of view.
III.1 Pulse-train solutions
Here we show for the first time that Eq. (7) allows for exact pulse–train solutions which are composed of the product of Jacobi elliptic functions. We find that this type of pulse shape can exist for Eq. (1) in three different classes.
Figure 1: Intensity profiles of pulse-train solutions with parameters $\gamma=2$, $k=0.9$, $\xi_{0}=0$ (a) pulse-train solution (16) when $\alpha=-1,$ $\sigma=-0.25,$ $\epsilon=-0.25$ (b) pulse-train solution (25) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5$ (c) pulse-train solution (34) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5.$
Figure 2: Intensity profiles of pulse-train solutions with parameter $k=0.999$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34). Other parameters are same as given in Fig. 1.
Figure 3: Evolution of pulse-train solutions with parameter $k=0.9$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34). Other parameters are same as given in Fig. 1.
Figure 4: Evolution of pulse-train solutions with parameter $k=0.999$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34). Other parameters are same as given in Fig. 1.
Figure 5: Dependence of the velocity of pulse-trains $v$ on the group velocity dispersion parameter $\alpha$ for $\sigma=1$ and different values of $\epsilon$, $\epsilon=-0.25$ (thick line), $\epsilon=-0.3$ (dashed line) and $\epsilon=-0.35$ (dotted line).
1. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{dn}\left(x,k\right)$-type
We introduce an appropriate ansatz for the propagating waveform solution of Eq. (7) as,
$\rho(\xi)=P\,\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)\mathrm{dn}\left(w(\xi-\xi_{0}),k\right),$ (9)
where $\xi_{0}$ is an arbitrary constant while $\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ and $\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)$ are the cosine and third kind of Jacobian elliptic functions of modulus $k$ taking values $10<k<$ 1.
Substituting the ansatz (9) into Eq. (7) and setting the coefficients of independent terms equal to zero, we obtain the following algebraic equations:
$\epsilon w^{4}(k^{4}-46k^{2}+61)-bw^{2}(k^{2}-5)=c,$ (10)
$60\epsilon w^{4}k^{2}(k^{2}-3)-6bw^{2}k^{2}+\gamma(k^{2}-1)P^{2}=0,$ (11)
$0120\epsilon w^{4}k^{2}+\gamma P^{2}=$ 0 . (12)
By solving the parametric equations (11) and (12), we get
$w=\frac{1}{4}\sqrt{\frac{8\alpha\epsilon-3\sigma^{2}}{5\epsilon^{2}(k^{2}+1)}},~{}~{}~{}~{}P=\pm\frac{k}{k^{2}+1}\sqrt{\frac{-6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (13)
with $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$. Further substitution of Eqs. (13) into Eq. (10) yields the parameter $c$ as
$c=\frac{(11k^{4}-86k^{2}+11)}{100\epsilon(k^{2}+1)^{2}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right)^{2},$ (14)
The wave number $\kappa$ follows from Eqs. (8) and (14) as
$\kappa=\frac{(11k^{4}-86k^{2}+11)}{100\epsilon^{3}(k^{2}+1)^{2}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (15)
Incorporating these results into Eq. (2), we obtain the following periodic wave solution to the extended NLSE (1):
$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{dn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (16)
where $\zeta=\tau-v^{-1}z-\xi_{0}$, with $\xi_{0}$ being the position of the wave at $z=0$. Note that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by the relations (5) and (6), respectively.
Physically, Eq. (16) describes a periodic wave with amplitude $P$ and inverse temporal width $w$ depending on all order of dispersion as well as nonlinearity.
We now introduce the quantity $\tau_{0}=1/w$ which represents the duration of the pulse [35]. Hence, we find that the relation between the pulse amplitude and duration is given by
$P\tau_{0}^{2}=-2k\left(-\Omega\right)^{1/2}.$ (17)
where the parameter $\Omega=30\epsilon/\gamma$ has been introduced for brevity.
2. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
We now make an ansatz for Eq. (7) as,
$\rho(\xi)=P\,\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right),$ (18)
where $\mathit{\mathrm{sn}}$ is Jacobi elliptic sine function of modulus $k$ which is in the range $10<k<$ 1.
Inserting this ansatz into Eq. (7) and equating the coefficients of independent terms, one obtains
$\epsilon w^{4}(61k^{4}-76k^{2}+16)+bw^{2}(5k^{2}-4)=c,$ (19)
$60\epsilon w^{4}k^{2}(3k^{2}-2)+6bw^{2}k^{2}+\gamma P^{2}=0,$ (20)
$0120\epsilon\eta^{4}k^{2}-\gamma P^{2}=$ 0 . (21)
Thus we have the following relations for the pulse inverse width $w$ and amplitude $\lambda$,
$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(4-3k^{2})}},~{}~{}~{}~{}P=\pm\frac{k}{4-3k^{2}}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (22)
with $\gamma\epsilon>0$ and $3\sigma^{2}>8\alpha\epsilon$. Additionally, Eqs. (18) and (21) yield the parameter $c$as
$c=\frac{b^{2}(244k^{2}-89k^{4}-144)}{100\epsilon(4-3k^{2})^{2}},$ (23)
Hence it follows from Eqs. (8) and (22) that the wave number $\kappa$ in the periodic wave solution is given by
$\kappa=\frac{(244k^{2}-89k^{4}-144)}{100\epsilon^{3}(4-3k^{2})^{2}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (24)
These equations yield the periodic wave solution as
$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (25)
where $\xi=\tau-v^{-1}z-\xi_{0}$. Noting that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by Eqs. (5) and (6), respectively.
For this solution, we find that the relation between the pulse amplitude and duration takes the form,
$P\tau_{0}^{2}=2k\left(\Omega\right)^{1/2}.$ (26)
3. Pulse-trains of the $\mathrm{dn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
Next we make an ansatz for Eq. (7) as,
$\rho(\xi)=P\,\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right),$ (27)
Inserting this ansatz into Eq. (7) and equating the coefficients of independent terms, one obtains
$\epsilon w^{4}(16k^{4}-76k^{2}+61)+bw^{2}(5k^{2}-4)=c,$ (28)
$60\epsilon w^{4}k^{2}(2k^{2}-3)-6bw^{2}k^{2}+\gamma P^{2}=0,$ (29)
$0120\epsilon w^{4}k^{2}-\gamma P^{2}=$ 0 . (30)
Thus we have the following relations for the pulse inverse width $\mu$ and amplitude $R$,
$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(2k^{2}-1)}},~{}~{}~{}~{}P=\pm\frac{k}{2k^{2}-1}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (31)
with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$ when $11/\sqrt{2}<k<$ 1 and for $8\alpha\epsilon>3\sigma^{2}$ when $0<k<1/\sqrt{2}$.
Moreover, Eqs. (18) and (21) yield the parameter $c$as
$c=\frac{b^{2}(64k^{2}-64k^{4}+11)}{100\epsilon(2k^{2}-1)^{2}},$ (32)
Hence it follows from Eqs. (8) and (22) that the wave number $\kappa$ in the periodic wave solution is given by
$\kappa=\frac{(64k^{2}-64k^{4}+11)}{100\epsilon^{3}(2k^{2}-1)^{2}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (33)
These equations yield the periodic wave solution as
$\psi(z,\tau)=P\,\mathrm{dn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (34)
where $\xi=\tau-v^{-1}z-\xi_{0}$. Note that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by the relations (5) and (6), respectively.
For this solution, we find that the relation between the pulse amplitude and duration takes the form,
$P\tau_{0}^{2}=2k\left(\Omega\right)^{1/2}.$ (35)
On comparing Eqs. (16), (25) and (34), we infer that under the influence of second-, third-, and fourth-order dispersions and Kerr nonlinearity the fiber medium allows the propagation of distinct types of pulse-trains which present a relation between the pulse amplitude and duration determined by the sign of the joint parameter $\Omega=30\epsilon/\gamma$ only. To our knowledge, the pulse-trains obtained above for the extended NLSE (1) are firstly reported in this paper.
Figure 1(a) depicts the intensity profile of the pulse-train solutions (16) of Eq. (1) for the parameters values: $\alpha=-1,$ $\gamma=2,$ $\sigma=-0.25,$ and $\epsilon=-0.25.$ The results for the pulse-train solutions (25) and (34) are illustrated in Figs. 1(b) and 1(c) for the values $\alpha=0.49,$ $\gamma=2,$ $\sigma=1,$ and $\epsilon=0.5$. Here the position $\xi_{0}$ of the waveform solutions at $z=0$ is choosen to be equal to zero. Also, the value of the elliptic modulus $k$ is taken as $k=0.9$. In Fig. 2, we show the evolution of these pulse-train solutions in the fiber medium. From these figures, we can see that all these pulse-trains are positive and exhibit an oscillating character. In addition, we can see that the pulse-train solutions (25) and (34)  present two peaks in each periodic unit, markedly different from the pulse train solution (16) which has a one peak structure.
It is also interesting to see from Figs. (3) and (4) that the shape of the pulse-train solution (25) is similar to that for the pulse-train solution (34) for the value of the elliptic modulus $k=0.999$.
The key observation is that the amplitudes, widths and wavenumbers of the derived pulse-train solutions are different whereas their velocity is the same. More importantly, the relation (6) shows that the velocity of the pulse-trains can be significantly reduced for suitable parameters of the waveguiding media, which may find application in developing slow light systems. In Fig. 5, we present the dependence of velocity on the group velocity dispersion parameter $\alpha$ of the fiber for three different values of the quartic dispersion parameter $\epsilon$. It can clearly be seen that the velocity decreases intensely in the magnitude with the increase of dispersion parameter $\alpha$. The decrease of velocity implies that the slowing motion of pulse-trains is possible in the transmitting medium when considering large dispersion parameter $\alpha$. It is also observed that the pulse velocity has large magnitude for small negative values of $\epsilon$. Hence, one concludes that the velocity of pulse-trains can be controlled by judicious choice of the dispersion parameters.
III.2 Quartic and dipole soliton solutions
We now focus on the long-wave limit of the new pulse-trains obtained above corresponding to the modulus $k=1$. In such a case, the periodic solution (16) degenerates to the quartic soliton solution [6]:
$\psi(z,\tau)=P_{0}~{}\mathrm{sech}^{2}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[i(\kappa z-\delta\tau+\theta)],$ (36)
where the inverse temporal width $w_{0}$, amplitude $P_{0}$ and wave number $\kappa$ of the soliton pulse are given by
$w_{0}=\frac{1}{4}\sqrt{\frac{8\alpha\epsilon-3\sigma^{2}}{10\epsilon^{2}}},$ (37)
$P_{0}=\pm\sqrt{\frac{-3}{10\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (38)
$\kappa=-\frac{4}{25\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (39)
with $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$. Also the velocity $v$ and frequency shift $\delta$ in this quartic soliton solution are defined by Eqs. (5) and (6).
In addion, when $k\rightarrow 1$, the periodic solution (25) transforms into our dipole soliton solution [7]:
$\psi(z,\tau)=P_{0}~{}\mathrm{sech}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\mathrm{th}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[i(\kappa z-\delta\tau+\theta)],$ (40)
where the inverse temporal width $w_{0}$, amplitude $P_{0}$ and wave number $\kappa$ of the soliton pulse take the form
$w_{0}=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}}},$ (41)
$P_{0}=\pm\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (42)
$\kappa=\frac{11}{100\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (43)
with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$. The velocity $v$ and frequency shift $\delta$ in this dipole soliton solution are given by Eqs. (5) and (6).
It is interesting to note that if one considers the limiting case with $k=1$, the periodic wave solution (34) also degenerates to the same soliton-like solutions (40).
IV Conclusion
In conclusion, we have discovered three classes of pulse-trains whose amplitudes, widths and wavenumbers are different with the distinctive property of exhibiting a relation between the amplitude and duration determined by the sign of a joint parameter uniquely. We have found that these structures have the same velocity, which depends only on the three dispersion parameters. It is demonstrated that higher-order dispersion may lead to the formation of double-humped pulse-trains besides the ones taking the usual form with single peak in each periodic unit. It is also shown that the pulse-trains are obtained without necessarily assuming a specified condition on the fiber parameters. The results presented here may be useful for experimental realization of pulse-trains in optical fibers and may help in stimulating more research in understanding their optical transmission properties.

### v2

@@
-Pulse-train propagation in nonlinear Kerr media governed by higher-order dispersion
+Pulse-train propagation in a nonlinear Kerr medium governed by higher-order dispersion
 Abstract
-We discover three novel classes of pulse-train waveforms in an optical Kerr nonlinear medium possessing all orders of dispersion up to the fourth order.
+We discover three novel classes of pulse-trains in an optical Kerr nonlinear medium possessing all orders of dispersion up to the fourth order.
 We show that both single- and double-humped pulse-trains can be formed in the nonlinear medium.
@@
 More importantly, we find that the relation between the amplitude and duration of all the newly obtained pulse-trains is determined by the sign of a joint parameter solely.
-The results show that those optical waves are general, in the sense that no specified conditions on the material parameters are assumed.
+The results show that those optical waves are general, in the sense that no specified conditions on the material parameters are taken on.
 Considering the long-wave limit, the derived pulse-trains degenerate to soliton pulses of the quartic and dipole kinds.
-pacs:
-05.45.Yv, 42.65.Tg
+Results in this study may be useful for experimental realization of pulse-trains in highly dispersive optical fibers and further understanding of their optical transmission properties.
+Keywords: Pulse-train, nonlinear Kerr medium, Higher-order dispersion, Quartic and dipole solitons, Nonlinear Schrödinger equation.
 I Introduction
-The transmission of soliton pulses in nonlinear media has been the focus of continuing interest, stimulated by their important potential applications in long-distance optical communication systems and optical switching devises [1].
-The formation of such wave packets in a monomode optical fiber results from the exact balance between the pulse-spreading effects of group-velocity dispersion and self-phase modulation effects due to the intensity-dependent refractive index of the fiber [2].
-Experimental observation of solitons have been reported in various optical systems, including single-mode optical fibers [3], femtosecond lasers [4] and bulk optical materials [5].
-In the picosecond regime, the propagation of light pulses inside a single-mode optical fiber was shown to be governed by the nonlinear Schrödinger equation (NLSE), which is commonly derived with use of the slowly varying envelope approximation [8].
-This physically relevant model allows for either bright or dark solitons in the anomalous and normal dispersion regimes, respectively [9].
-However, for much shorter light pulses, there is also appearance of additional higher-order effects in the optical system.
-For example, the third-order dispersion significantly influence the short pulses whose widths are nearly 50 femtosecond as they propagate through the nonlinear medium [10, 11].
-The fourth-order dispersion is also found to be important when pulses are shorter than 10 femtosecond [10, 11].
-In such a situation, extensions of the NLSE model that take into account the influence of diverse physical phenomena on short-pulse propagation and generation are therefore essential to describe the wave dynamics accurately.
-One notes that higher-order effects introduce several new important phenomena into the system dynamics that are absent in Kerr nonlinear medium.
-A great interest has been given recently to the generation of pulse-trains in various physical systems, including birefringent optical fibers [12, 13, 14, 15, 16], negative-index materials [17], Bose-Einstein condensates [18], etc.
-It can be relevant to note that such kind of pulse shape is interesting in itself as another type of physically relevant nonlinear waves in optical fibers besides the soliton-type waveform.
+The transmission of envelope solitons in waveguiding media has been the focus of continuing interest, stimulated by their important potential applications in optical communication systems and optical switching devises [1].
+The formation of such wave packets in a monomode optical fiber results from an interplay between nonlinear self-phase modulation and group-velocity dispersion effects [2].
+One notes that the experimental observation of solitons has been demonstrated in diverse optical systems, including nonlinear optical fibers [3], glass waveguides [4] and femtosecond ring lasers [5].
+Recently, the properties of soliton pulses have been studied extensively due to their important applications for optical transmission and information processing [6].
+Usually, the evolution of picosecond light pulses inside a monomode fiber has been more generally governed by the nonlinear Schrödinger equation (NLSE) that is commonly derived with use of the slowly varying envelope approximation [7].
+Such a pertinent model has admitted the bright pulse soliton solution in the anomalous dispersion regime while in the normal dispersion region of the fiber it has exhibited the dark-type soliton solution [8].
+However, when much shorter optical pulses have to be injected in the fiber, many higher-order dispersive and nonlinear effects have become significant and should have been taken into consideration [9].
+For example, the third-order dispersion has considerably influenced the short pulses whose widths are about $50$ femtoseconds as they propagate through the nonlinear medium [10, 11].
+The quartic dispersion has been also found to be significant when the light pulses are shorter than $10$ femtoseconds [10, 11].
+In such a case, extensions of the cubic model taking into account the specific contributions of various higher-order effects on the light propagation have been therefore essential to describe the wave dynamics accurately [12].
+It has been worth to noting that the occurrence of higher-order processes gives rise to a multitude of novel phenomena which are not found in the case of the pure Kerr media [13].
+One notes that besides the NLSE and its variants, important results have been also obtained with recent studies discussing diverse nonlinear models which find applications in various sciences and engineering fields, including magnetic field [14], fluid dynamics [15, 16], magneto-optics [17], quantum mechanics [18], and plasma physics [19].
+These models have been used to understand and simulate complex phenomena like solitons in magnetic materials, shallow-water waves in oceans, and the propagation of nonlinear waves in various media.
+We also note that significant results have been recently obtained by studying the existence and dynamics of various new soliton structures in nonlinear physical systems governed by the NLS-type equations, thus illustrating their significance in fields like optics and Bose-Einstein condensates [20, 21, 22].
+Techniques like bilinear auto-Bäcklund transformations and similarity reductions have been found to be crucial for finding solutions for such models and understanding system behaviors.
+Nowadays, a great interest has been given to the generation of pulse-trains in various physical systems, including birefringent optical fibers [23, 24, 25, 26, 27], negative-index materials [28], Bose-Einstein condensates [29], etc.
+It can be relevant to note that such type of pulse shape is attracting in itself as another kind of physically relevant nonlinear waves in optical fibers besides the soliton waveforms.
 These structures exhibit many interesting properties that make them potentially in problems of signal propagation in nonlinear fibers [30].
-It is intersting to note that pulse-trains can be produced experimentally.
-For example, the formation of cnoidal wave has been recently demonstrated in a fiber laser either in the net normal or net anomalous cavity dispersion regime [20].
-Importantly, such arrays of pulses may be created by the modulational instability of the continuous-wave signal [21], or from a dual-frequency pump [22].
-In optics, such nonlinear waves play a pivotal role in the analysis of the data transmission in fiber-optic telecommunications links [23].
+It is worth mentioning that pulse-trains can be produced experimentally.
+For example, the formation of cnoidal waves has been recently demonstrated in a fiber laser either in the net anomalous or net normal cavity dispersion regime [31].
+Importantly, such pulse-trains may be created from a dual-frequency pump [32] or by the modulational instability of the continuous-wave signal [33].
+In optics, such nonlinear structures play a pivotal role in analyzing the data transmission in optical communication links [34].
+Recently, a kind of extended NLSE with second-, third-, and fourth-order dispersion terms has been shown to support a stable solitary wave solution with a $\mathrm{sech}^{2}$ shape which is also called quartic soliton [35].
+Such kind of stable localized pulses could find practical applications in communications, ultrafast lasers, and slow-light devices [35].
+Due to its physical importance in describing femtosecond pulse propagation in highly dispersive optical fibers, such equation has been analyzed from different points of view [35, 36, 37].
+With regards to the periodic waves, one of the present authors [37] has analytically solved this model and obtained results for periodic wave solutions of the $\mathrm{cn}^{2}$ type.
+However, analytical pulse-train solutions of this extended NLSE that are expressed in terms of products of two different Jacobi elliptic functions have not been reported yet.
 Identification of new kinds of pulse-trains in optical systems is an issue worth considering.
 It is pertinent to mention that the finding of such structures may lead to novel dynamical behaviors of waves and opens the possibility for experimental investigations of various wave phenomena.
-In this work, we demonstrate for the first time that an optical Kerr nonlinear medium exhibiting all orders of dispersion up to the fourth order supports the existence of three novel classes of pulse-trains which take either single- or double-hump forms.
-The paper is structured as follows.
-Section II provides the analytic framework necessary to find the traveling wave solutions of the extended NLSE model that governs the transmission of extremely short pulses throught an optical fiber medium exhibiting all orders of dispersion up to the fourth order.
-In Section III, we derive analytical pulse-train solutions of the model and their characteristics.
+In this study, we show for the first time that an optical Kerr nonlinear system displaying dispersion effects up to the fourth order supports the existence of three novel classes of pulse-trains which take either single- or double-hump forms.
+The main novelty of these solutions is that the periodic pulse-trains that are composed of products of Jacobian elliptic functions are firstly reported for the extended NLSE with third- and fourth-order dispersions, the interesting amplitude-duration relationship for these structures is found, and the propagation of pulse-train solutions is firstly represented by numerical simulations.
+In particular, the outcomes presented below constitute the first analytical demonstration of possible pulse-train generation in highly dispersive optical fibers, which may have potential application for the further experiments and research in nonlinear optics.
+These results will be helpful to nonlinear optics since this kind of nonlinear waves serves as a model of pulse train propagation in optics fibers [34] due to their structural stability with respect to the small input profile perturbations and collisions [38].
+The rest of paper is structured as follows.
+In Sec.
+II, we present the analytic framework necessary to find the traveling wave solutions of the extended NLSE model which governs the transmission of extremely short light pulses through a fiber system depicting higher-order dispersion effects.
+In Sec.
+III, we derive three different analytic pulse-train solutions of the governing equation and their characteristics.
 We also discuss here the propagation properties of obtained pulse-trains in the fiber medium.
-Finally, some conclusions are given in Sec.
+In Sec.
+IV, we numerically simulate the propagation of the obtained pulse-train solutions.
+The stability of the solutions is discussed numerically in Sec.
+V.
+The conclusions of this paper will be then summarized in Sec.
 VI.
 II Model and traveling waves
-We consider the evolution of an optical pulse in a highly dispersive fiber medium having a Kerr nonlinearity.
-The governing model for the slowly varying envelope $\psi$ of the optical pulse can be described by the extended NLSE [6, 7],
-$i\frac{\partial\psi}{\partial z}=\alpha\frac{\partial\psi^{2}}{\partial\tau^{2}}+i\sigma\frac{\partial\psi^{3}}{\partial\tau^{3}}-\epsilon\frac{\partial\psi^{4}}{\partial\tau^{4}}-\gamma\left|\psi\right|^{2}\psi.$ (1)
-where $\tau=t-\beta_{1}z$ is the retarded time, $z$ is the longitudinal coordinate, $\alpha=\beta_{2}/2$, $\sigma=\beta_{3}/6$, $\epsilon=\beta_{4}/24$, and $\gamma$ is the nonlinear parameter.
-Here the parameters $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ are the first-, second-, third-, and fourth-order dispersion coefficients, respectively.
+We consider the propagation of ultrashort light pulses through a nonlinear fiber medium exhibiting the effects of higher-order dispersions as well as self-phase modulation nonlinearity.
+The governing model for the slowly varying envelope $\psi$ of the light pulse can be described by the extended NLSE [35, 36],
+$\mathrm{i}\frac{\partial\psi}{\partial z}=\alpha\frac{\partial^{2}\psi}{\partial\tau^{2}}+\mathrm{i}\text{\/}\sigma\frac{\partial^{3}\psi}{\partial\tau^{3}}-\epsilon\frac{\partial^{4}\psi}{\partial\tau^{4}}-\gamma\left|\psi\right|^{2}\psi,$ (1)
+where $z$ is the longitudinal coordinate, $\tau=t-\beta_{1}z$ is the retarded time, $\alpha=\beta_{2}/2$, $\sigma=\beta_{3}/6$, $\epsilon=\beta_{4}/24$, while $\gamma$ is the fiber nonlinearity coefficient that results in self-phase modulation.
+Here $\beta_{j}$ ($j=1,2,3,4$) represent, respectively, the dispersion coefficients of the first, second, third, and fourth order.
+As an important generalization, the extended NLSE (1) includes, besides the group-velocity dispersion and self-phase modulation terms that constitute the cubic NLS model, the effects of third- and fourth-order dispersions.
+These higher-order effects arise in highly dispersive optical systems including dispersion-shifted fibers [39], photonic crystal waveguides [40], and silicon-based waveguides [41].
+It is relevant to note here that the third- and fourth-order dispersions occur in the region of minimum group-velocity dispersion [42], in which they become crucial for describing femtosecond pulse behavior.
 Theoretical investigations based on Eq.
 (1) are mainly concentrated on the existence and propagation properties of localized pulses due to their fundamental importance in the understanding of various physical phenomena in the system.
-For example, Kruglov and Harvey [6] demonstrated the existence of stable solitary waves with a $\mathrm{sech}^{2}$ shape for the model.
-Triki and Kruglov [7] discussed the dipole soliton solution of Eq.
+As previously mentioned, Kruglov and Harvey [35] demonstrated the existence of stable solitary waves with a $\mathrm{sech}^{2}$ shape for this model.
+Triki and Kruglov [36] discussed the dynamics of dipole soliton waveforms of Eq.
 (1) in the presence the inhomogeneities of media.
+Kruglov [37] obtained the periodic wave solutions that take the $\mathrm{cn}^{2}$ shape for this underlying equation [37].
 Cavalcanti et al.
-[24] analyzed the modulational instability phenomena of Eq.
-(1) in the region of the minimum group-velocity dispersion.
+[42] analyzed the modulation instability of the model (1) in the region near the zero-dispersion wavelength.
 Karpman et al.
-[25, 26] investigated the time behavior of the amplitudes, velocities, and other parameters of radiating solitons.
-Shagalov [27] studied the effect of the third- and fourth-order dispersions on the modulational instability.
+[43, 44] investigated the resonant radiation and evolution of a soliton described by the model (1).
+Shagalov [45] studied the influence of high dispersion terms in the governing equation (1) on the modulational instability of nonlinear waves.
 Roy et al.
-[29] discussed the role of third- and fourth-order dispersions in the radiation emitted by fundamental soliton pulses in the form of dispersive waves within the framework of the dimensionless form of the model (1).
-Kruglov [28] obtained the periodic wave solutions that take the $\mathrm{cn}^{2}$ shape for this governing equation [28].
-In what follows, we present three novel exact pulse-train solutions with interesting properties for the extended NLSE (1), which are obtained without necessarily assuming a specified condition on the fiber parameters.
-In order to find some interesting analytical solutions of the model (1), we consider a solution of the form
-$\psi(z,\tau)=\rho(\xi)\exp[i(\kappa z-\delta\tau+\theta)],$ (2)
-where $\rho(\xi)$ is a real function, $\xi=\tau-qz$ is a traveling coordinate with $q=v^{-1}$ being the inverse velocity.
-Also, $\kappa$ and $\delta$ are the respective real parameters representing the wavenumber and frequency shift, while $\theta$ represents the phase of the pulse at $z=0$.
-Substitution of Eq. (1) into the extended NLSE (2) leads to a set of coupled differential equations:
-$(\sigma+4\epsilon\delta)\frac{d^{3}\rho}{d\xi^{3}}+(q-2\alpha\delta-3\sigma\delta^{2}-4\epsilon\delta^{3})\frac{d\rho}{d\xi}=0,$ (3)
-$\epsilon\frac{d^{4}\rho}{d\xi^{4}}-(\alpha+3\sigma\delta+6\epsilon\delta^{2})\frac{d^{2}\rho}{d\xi^{2}}+\gamma\rho^{3}-(\kappa-\alpha\delta^{2}-\sigma\delta^{3}-\epsilon\delta^{4})\rho=0.$ (4)
-Here, we refer the solution of the extended NLSE where the function $\psi(z,\tau)$ is given by Eq. (2) with $\rho(\xi)\neq\mathrm{constant}$ as non-plain wave or traveling wave solution.
-The system of Eqs. (3) and (4) with $\epsilon\neq 0$ yields the non-plain wave solutions if and only if the next relations are satisfied:
+[46] discussed the roles of high-order dispersions in the generation and control of dispersive waves.
+In what follows, we present three novel analytic pulse-train solutions with interesting properties for the extended NLSE (1), which are obtained without necessarily assuming a specified condition on the fiber parameters.
+Here, the ansatz method which is one of the effective and powerful techniques for finding analytic localized and periodic wave solutions of nonlinear evolution equations [14, 15, 16], is used to obtain various types of pulse-trains solutions for the studied model.
+In order to find some interesting analytical solutions of the model (1), we make the transformation [35, 36],
+$\psi(z,\tau)=\rho(\xi)\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (2)
+with $\rho(\xi)$ is the real amplitude and $\xi=\tau-qz$ is the traveling coordinate.
+Here the inverse velocity $q=v^{-1}$, the frequency shift $\delta$, the wave number $\kappa$, and the phase constant $\theta$ (at $z=0$) are all real parameters to be determined.
+Inserting Eq. (2) into the extended NLSE (1), we get the differential equations:
+$(\sigma+4\epsilon\delta)\frac{\mathrm{d}^{3}\rho}{\mathrm{d}\xi^{3}}+(q-2\alpha\delta-3\sigma\delta^{2}-4\epsilon\delta^{3})\frac{\mathrm{d}\rho}{\mathrm{d}\xi}=0,$ (3)
+$\epsilon\frac{\mathrm{d}^{4}\rho}{\mathrm{d}\xi^{4}}-(\alpha+3\sigma\delta+6\epsilon\delta^{2})\frac{\mathrm{d}^{2}\rho}{\mathrm{d}\xi^{2}}+\gamma\rho^{3}-(\kappa-\alpha\delta^{2}-\sigma\delta^{3}-\epsilon\delta^{4})\rho=0.$ (4)
+Here, one refers the waveform solution of the model under consideration where the envelope function $\psi(z,\tau)$ is defined by the representation (2) with $\rho(\xi)\neq\mathrm{constant}$ as traveling wave or non-plain wave solution.
+Equations (3) and (4) with non-zero quartic dispersion coefficient (i.e., $\epsilon\neq 0$) give the non-plain wave solutions when the following relations are fulfilled:
 $q=2\alpha\delta+3\sigma\delta^{2}+4\epsilon\delta^{3},~~~~\delta=-\frac{\sigma}{4\epsilon}.$ (5)
-Note that Eq.
-(3) is satisfied for an arbitrary function $\rho(\xi)$ according to the conditions in Eq.
+Notice that the differential equation (3) is fulfilled for an arbitrary amplitude function $\rho(\xi)$ in accordance with the conditions in Eq.
 (5) with $\epsilon\neq 0$.
-The relations in Eq.
-(5) leads to the next expression for the velocity $v=1/q$ defined in the retarded frame,
+Additionally, Eq.
+(5) allows us to obtain the following expression for the wave velocity $v=1/q$,
 $v=\frac{8\epsilon^{2}}{\sigma(\sigma^{2}-4\alpha\epsilon)}.$ (6)
 The latter relation shows that the dispersion parameters $\beta_{i}$ $(i=2,3,4)$ included in the parameters $\alpha$, $\sigma$ and $\epsilon$ have a major influence on the wave velocity.
-This implies that the velocity of the pulse can be controlled by an appropriate manipulation of the dispersion parameters.
-With use of the relations in Eq. (5), the system of Eqs. (3) and (4) reduces to the ordinary nonlinear differential equation,
-$\epsilon\frac{d^{4}\rho}{d\xi^{4}}+b\frac{d^{2}\rho}{d\xi^{2}}-c\rho+\gamma\rho^{3}=0,~{}~{}~{}~{}~{}~{}~{}$ (7)
-where the parameters $b$ and $c$ are
+This implies that the velocity of the pulse can be efficiently controlled by varying these parameters.
+Now using Eqs. (4) and (5), one can find an equation for $\rho$ as
+$\epsilon\frac{\mathrm{d}^{4}\rho}{\mathrm{d}\xi^{4}}+b\frac{\mathrm{d}^{2}\rho}{\mathrm{d}\xi^{2}}-c\rho+\gamma\rho^{3}=0,$~~~~~~~ (7)
+with
 $b=\frac{3\sigma^{2}}{8\epsilon}-\alpha,~~~~c=\kappa+\frac{\sigma^{2}}{16\epsilon^{2}}\left(\frac{3\sigma^{2}}{16\epsilon}-\alpha\right).$ (8)
-Equation (7) describes the evolution of the field amplitude in the highly dispersive optical fiber.
-Now the natural question is whether the amplitude equation (7) possesses exact pulse-train solutions without any specified condition on the coefficients or not.
-In following, we have identified novel families of pulse-train solutions with some interesting properties for this evolutional equation in presence of all physical parameters.
+Equation (7) describes the evolution of the wave amplitude in the highly dispersive nonlinear fiber.
+Now the natural question is whether the amplitude equation (7) possesses analytic pulse-train solutions without any specified condition on the coefficients or not.
+In following, we have identified novel families of pulse-train solutions with some interesting properties for this evolution equation taking into account the influence of all linear and nonlinear processes on pulse propagation.
 III Results and discussion
@@
 Here we show for the first time that Eq.
-(7) allows for exact pulse–train solutions which are composed of the product of Jacobi elliptic functions.
+(7) allows for analytic pulse-train solutions which are expressed in terms of the product of different elliptic functions.
 We find that this type of pulse shape can exist for Eq.
 (1) in three different classes.
-Figure 1: Intensity profiles of pulse-train solutions with parameters $\gamma=2$, $k=0.9$, $\xi_{0}=0$ (a) pulse-train solution (16) when $\alpha=-1,$ $\sigma=-0.25,$ $\epsilon=-0.25$ (b) pulse-train solution (25) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5$ (c) pulse-train solution (34) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5.$
-Figure 2: Intensity profiles of pulse-train solutions with parameter $k=0.999$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34).
-Other parameters are same as given in Fig.
+Figure 1: Intensity profiles of pulse-train solutions with parameters $\gamma=2$, $k=0.9$, $\xi_{0}=0$; (a) pulse-train solution (16) when $\alpha=-1,$ $\sigma=-0.25,$ $\epsilon=-0.25$; (b) pulse-train solution (25) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5$; (c) pulse-train solution (34) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5.$
+Figure 2: Intensity profiles of pulse-train solutions with parameter $k=0.9$; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34).
+Other parameters are the same as given in Fig.
 1.
-Figure 3: Evolution of pulse-train solutions with parameter $k=0.9$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34).
-Other parameters are same as given in Fig.
+Figure 3: Evolution of pulse-train solutions with parameter $k=0.999$; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34).
+Other parameters are the same as given in Fig.
 1.
-Figure 4: Evolution of pulse-train solutions with parameter $k=0.999$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34).
-Other parameters are same as given in Fig.
+Figure 4: Evolution of pulse-train solutions with parameter $k=0.999$; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34).
+Other parameters are the same as given in Fig.
 1.
-Figure 5: Dependence of the velocity of pulse-trains $v$ on the group velocity dispersion parameter $\alpha$ for $\sigma=1$ and different values of $\epsilon$, $\epsilon=-0.25$ (thick line), $\epsilon=-0.3$ (dashed line) and $\epsilon=-0.35$ (dotted line).
+Figure 5: Dependence of the velocity of pulse-trains $v$ on the group velocity dispersion parameter $\alpha$ for $\sigma=1$ and different values of $\epsilon$; $\epsilon=-0.25$ (thick line), $\epsilon=-0.3$ (dashed line) and $\epsilon=-0.35$ (dotted line).
+Figure 6: Evolution of the intensity distribution of the extended NLSE; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34). Parameters are given in the text.
 1. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{dn}\left(x,k\right)$-type
-We introduce an appropriate ansatz for the propagating waveform solution of Eq. (7) as,
-$\rho(\xi)=P\,\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)\mathrm{dn}\left(w(\xi-\xi_{0}),k\right),$ (9)
-where $\xi_{0}$ is an arbitrary constant while $\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ and $\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)$ are the cosine and third kind of Jacobian elliptic functions of modulus $k$ taking values $10<k<$ 1.
-Substituting the ansatz (9) into Eq. (7) and setting the coefficients of independent terms equal to zero, we obtain the following algebraic equations:
+We introduce an appropriate ansatz for the propagating waveform solution of Eq. (7) as
+$\rho(\xi)=P\,\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)\mathrm{cn}\left(w(\xi-\xi_{0}),k\right),$ (9)
+with cosine and the third kind of Jacobian elliptic functions $\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ and $\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)$ of the modulus number $k$ ($0<k<1$).
+Here $P$ is the amplitude of the wave to be determined and $\xi_{0}$ is its initial position (at $z=0$).
+Substituting the ansatz (9) into Eq.
+(7) and setting the coefficients of $\mathrm{dn}^{j}\left(w(\xi-\xi_{0}),k\right)\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ functions (where $j=1,3,5$) equal to zero, one gets the algebraic equations:
 $\epsilon w^{4}(k^{4}-46k^{2}+61)-bw^{2}(k^{2}-5)=c,$ (10)
 $60\epsilon w^{4}k^{2}(k^{2}-3)-6bw^{2}k^{2}+\gamma(k^{2}-1)P^{2}=0,$ (11)
-$0120\epsilon w^{4}k^{2}+\gamma P^{2}=$ 0 . (12)
+$120\epsilon w^{4}k^{2}+\gamma P^{2}=0.$ (12)
 By solving the parametric equations (11) and (12), we get
-$w=\frac{1}{4}\sqrt{\frac{8\alpha\epsilon-3\sigma^{2}}{5\epsilon^{2}(k^{2}+1)}},~{}~{}~{}~{}P=\pm\frac{k}{k^{2}+1}\sqrt{\frac{-6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (13)
+$w=\frac{1}{4}\sqrt{\frac{8\alpha\epsilon-3\sigma^{2}}{5\epsilon^{2}(k^{2}+1)}},~~~~P=\pm\frac{k}{k^{2}+1}\sqrt{\frac{-6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (13)
 with $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$. Further substitution of Eqs. (13) into Eq. (10) yields the parameter $c$ as
-$c=\frac{(11k^{4}-86k^{2}+11)}{100\epsilon(k^{2}+1)^{2}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right)^{2},$ (14)
-The wave number $\kappa$ follows from Eqs. (8) and (14) as
+$c=\frac{(11k^{4}-86k^{2}+11)}{100\epsilon(k^{2}+1)^{2}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right)^{2}.$ (14)
+Then using Eqs. (8) and (14), we can obtain the expression for the wave number $\kappa$:
 $\kappa=\frac{(11k^{4}-86k^{2}+11)}{100\epsilon^{3}(k^{2}+1)^{2}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (15)
-Incorporating these results into Eq. (2), we obtain the following periodic wave solution to the extended NLSE (1):
-$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{dn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (16)
-where $\zeta=\tau-v^{-1}z-\xi_{0}$, with $\xi_{0}$ being the position of the wave at $z=0$.
-Note that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by the relations (5) and (6), respectively.
-Physically, Eq. (16) describes a periodic wave with amplitude $P$ and inverse temporal width $w$ depending on all order of dispersion as well as nonlinearity.
+Based on these findings, we get a novel periodic wave solution for the model (1) as
+$\psi(z,\tau)=P\,\mathrm{dn}\left(w\zeta,k\right)\mathrm{cn}\left(w\zeta,k\right)\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (16)
+where $\zeta=\tau-v^{-1}z-\xi_{0}$. Note that the frequency shift $\delta$ and velocity $v$ of this periodic waveform are defined by the relations (5) and (6), respectively.
+From Eq.
+(13), one can see that the amplitude $P$ and inverse temporal width $w$ of this periodic wave crucially depend on various dispersion effects as well as nonlinearity.
+This result indicates that the contributions of dispersions and self-phase modulation processes are essential for this newly found pulse-train to be formed in the optical fiber medium.
 We now introduce the quantity $\tau_{0}=1/w$ which represents the duration of the pulse [47]. Hence, we find that the relation between the pulse amplitude and duration is given by
-$P\tau_{0}^{2}=-2k\left(-\Omega\right)^{1/2}.$ (17)
-where the parameter $\Omega=30\epsilon/\gamma$ has been introduced for brevity.
+$P\tau_{0}^{2}=-2k\left(-\Omega\right)^{1/2},$ (17)
+where the parameter $\Omega=30\epsilon/\gamma$ has been introduced for brevity. Note that here, we considered the ++ sign in all the expressions of the wave amplitude.
 2. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
-We now make an ansatz for Eq. (7) as,
+We now make an ansatz for Eq. (7) as
 $\rho(\xi)=P\,\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right),$ (18)
-where $\mathit{\mathrm{sn}}$ is Jacobi elliptic sine function of modulus $k$ which is in the range $10<k<$ 1.
-Inserting this ansatz into Eq. (7) and equating the coefficients of independent terms, one obtains
+with cosine and sine Jacobian elliptic functions $\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ and $\mathrm{sn}\left(w(\xi-\xi_{0}),k\right)$ of the modulus number $k$ ($0<k<1$).
+Upon substituting of this ansatz into Eq.
+(7) and then setting the coefficients of $\mathrm{cn}^{j}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right)$ functions (where $j=1,3,5$) equal to zero, we obtain the algebraic equations:
 $\epsilon w^{4}(61k^{4}-76k^{2}+16)+bw^{2}(5k^{2}-4)=c,$ (19)
 $60\epsilon w^{4}k^{2}(3k^{2}-2)+6bw^{2}k^{2}+\gamma P^{2}=0,$ (20)
-$0120\epsilon\eta^{4}k^{2}-\gamma P^{2}=$ 0 . (21)
-Thus we have the following relations for the pulse inverse width $w$ and amplitude $\lambda$,
-$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(4-3k^{2})}},~{}~{}~{}~{}P=\pm\frac{k}{4-3k^{2}}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (22)
-with $\gamma\epsilon>0$ and $3\sigma^{2}>8\alpha\epsilon$. Additionally, Eqs. (18) and (21) yield the parameter $c$as
-$c=\frac{b^{2}(244k^{2}-89k^{4}-144)}{100\epsilon(4-3k^{2})^{2}},$ (23)
-Hence it follows from Eqs. (8) and (22) that the wave number $\kappa$ in the periodic wave solution is given by
+$120\epsilon\eta^{4}k^{2}-\gamma P^{2}=0.$ (21)
+Based on these equations, one obtains the relations for the inverse width $w$ of the wave and its amplitude $P$ as
+$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(4-3k^{2})}},~~~~P=\pm\frac{k}{4-3k^{2}}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (22)
+with $\gamma\epsilon>0$ and $3\sigma^{2}>8\alpha\epsilon$. Additionally, by using Eqs. (19) and (22), we have the expression:
+$c=\frac{b^{2}(244k^{2}-89k^{4}-144)}{100\epsilon(4-3k^{2})^{2}}.$ (23)
+Then from Eqs. (8) and (23), we find that the wave number $\kappa$ of the periodic wave solution takes the form as
 $\kappa=\frac{(244k^{2}-89k^{4}-144)}{100\epsilon^{3}(4-3k^{2})^{2}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (24)
-These equations yield the periodic wave solution as
-$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (25)
-where $\xi=\tau-v^{-1}z-\xi_{0}$. Noting that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by Eqs. (5) and (6), respectively.
+These results yield a new kind of pulse-train solution for the NLSE model (1) as
+$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (25)
+where $\xi=\tau-v^{-1}z-\xi_{0}$. Noting that the frequency shift $\delta$ and velocity $v$ in this periodic structure are defined by the relations (5) and (6), respectively.
 For this solution, we find that the relation between the pulse amplitude and duration takes the form,
@@
 3. Pulse-trains of the $\mathrm{dn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
-Next we make an ansatz for Eq. (7) as,
-$\rho(\xi)=P\,\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right),$ (27)
-Inserting this ansatz into Eq. (7) and equating the coefficients of independent terms, one obtains
+Next we make an ansatz for Eq. (7) as
+$\rho(\xi)=P\,\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right).$ (27)
+Substituting this ansatz into the amplitude equation (7) and then setting the coefficients of $\mathrm{dn}^{j}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right)$ functions (where $j=1,3,5$) equal to zero, one finds the algebraic equations:
 $\epsilon w^{4}(16k^{4}-76k^{2}+61)+bw^{2}(5k^{2}-4)=c,$ (28)
 $60\epsilon w^{4}k^{2}(2k^{2}-3)-6bw^{2}k^{2}+\gamma P^{2}=0,$ (29)
-$0120\epsilon w^{4}k^{2}-\gamma P^{2}=$ 0 . (30)
-Thus we have the following relations for the pulse inverse width $\mu$ and amplitude $R$,
-$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(2k^{2}-1)}},~{}~{}~{}~{}P=\pm\frac{k}{2k^{2}-1}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (31)
-with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$ when $11/\sqrt{2}<k<$ 1 and for $8\alpha\epsilon>3\sigma^{2}$ when $0<k<1/\sqrt{2}$.
-Moreover, Eqs. (18) and (21) yield the parameter $c$as
-$c=\frac{b^{2}(64k^{2}-64k^{4}+11)}{100\epsilon(2k^{2}-1)^{2}},$ (32)
-Hence it follows from Eqs. (8) and (22) that the wave number $\kappa$ in the periodic wave solution is given by
+$120\epsilon w^{4}k^{2}-\gamma P^{2}=0.$ (30)
+Solving these equations, one can get the pulse parameters $w$ and $P$ as
+$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(2k^{2}-1)}},~~~~P=\pm\frac{k}{2k^{2}-1}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (31)
+with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$ when $1/\sqrt{2}<k<1$ and $8\alpha\epsilon>3\sigma^{2}$ when $0<k<1/\sqrt{2}$.
+Moreover, the use of Eqs. (28) and (31) lead to an expression for the parameter cc\ as
+$c=\frac{b^{2}(64k^{2}-64k^{4}+11)}{100\epsilon(2k^{2}-1)^{2}}.$ (32)
+Now using Eqs. (8) and (32), we can determine the wave number $\kappa$ as
 $\kappa=\frac{(64k^{2}-64k^{4}+11)}{100\epsilon^{3}(2k^{2}-1)^{2}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (33)
-These equations yield the periodic wave solution as
-$\psi(z,\tau)=P\,\mathrm{dn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (34)
-where $\xi=\tau-v^{-1}z-\xi_{0}$. Note that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by the relations (5) and (6), respectively.
+Thus, we find a novel pulse-train solution for the NLSE model (1) as
+$\psi(z,\tau)=P\,\mathrm{dn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (34)
+where $\xi=\tau-v^{-1}z-\xi_{0}$. Here $\delta$ and $v$ can be worked out by using the relations (5) and (6), respectively.
 For this solution, we find that the relation between the pulse amplitude and duration takes the form,
@@
 On comparing Eqs.
-(16), (25) and (34), we infer that under the influence of second-, third-, and fourth-order dispersions and Kerr nonlinearity the fiber medium allows the propagation of distinct types of pulse-trains which present a relation between the pulse amplitude and duration determined by the sign of the joint parameter $\Omega=30\epsilon/\gamma$ only.
-To our knowledge, the pulse-trains obtained above for the extended NLSE (1) are firstly reported in this paper.
-Figure 1(a) depicts the intensity profile of the pulse-train solutions (16) of Eq.
-(1) for the parameters values: $\alpha=-1,$ $\gamma=2,$ $\sigma=-0.25,$ and $\epsilon=-0.25.$ The results for the pulse-train solutions (25) and (34) are illustrated in Figs.
+(17), (26) and (35), we infer that under the effects of quadratic, cubic and quartic dispersions and self-phase modulation nonlinearity, the fiber system allows the propagation of distinct types of pulse-trains which present a relation between the pulse amplitude and duration determined by the sign of the joint parameter $\Omega=30\epsilon/\gamma$ only.
+To our knowledge, the pulse-trains derived above for Eq.
+(1) are firstly presented in this work.
+Figure 1(a) depicts the intensity profile of the pulse-train solution (16) of the model (1) for the parameters values: $\alpha=-1,$ $\gamma=2,$ $\sigma=-0.25,$ and $\epsilon=-0.25.$ The results for the pulse-train solutions (25) and (34) are shown in Figs.
 1(b) and 1(c) for the values $\alpha=0.49,$ $\gamma=2,$ $\sigma=1,$ and $\epsilon=0.5$.
-Here the position $\xi_{0}$ of the waveform solutions at $z=0$ is choosen to be equal to zero.
-Also, the value of the elliptic modulus $k$ is taken as $k=0.9$.
+One notes here that the value of initial position $\xi_{0}$ (at $z=0$) of the nonlinear waves is chosen zero.
+Also, the modulus number $k$ is selected to be $k=0.9$.
+The evolution of these pulse-train solutions in the fiber medium are displayed in Fig.
+2.
+From these figures, we observe that the three types of pulse-train solutions are positive and exhibit an oscillating behavior.
+In addition, we can see that the pulse-train solutions (25) and (34) present two peaks in each periodic unit, markedly different from the pulse train solution (16) which has a one peak structure.
+It is also interesting to see from Figs. (3) and (4) that the shape of the pulse-train solution (25) is similar to that for the pulse-train solution (34) for the modulus number $k=0.999$.
+A key observation is that the amplitudes, widths and wave numbers of the derived pulse-train solutions are different whereas their velocity is the same.
+More importantly, the expression (6) demonstrates that the velocity of these pulse-trains can be considerably diminished for suitable parameters of the waveguiding media, which have application value in the development of slow-light systems.
+Recently, Li and Huang [48] studied the modification of a slow-light soliton in three-level atomic systems.
+They derived a high-order NLSE incorporating third-order dispersion, delay in nonlinear refractive index, modulation induced by the microwave field, nonlinear dispersion, and correction terms of linear and differential absorptions; and demonstrated that an obvious decrease of propagating velocity of the soliton can be obtained in the presence of the microwave field.
+Note that photonic crystal waveguides (PhCWs) represent interesting slow light structures due to their confinement and versatile dispersion properties including slow light and possible control of group velocity dispersion effects [49, 50, 51].
+In order to strictly answer the question of possibility of applying the newly found pulse-trains in the development of slow-light systems much further research is needed.
+A detailed investigation of this problem will be reported in a forthcoming publication.
 In Fig.
-2, we show the evolution of these pulse-train solutions in the fiber medium.
-From these figures, we can see that all these pulse-trains are positive and exhibit an oscillating character.
-In addition, we can see that the pulse-train solutions (25) and (34)  present two peaks in each periodic unit, markedly different from the pulse train solution (16) which has a one peak structure.
-It is also interesting to see from Figs.
-(3) and (4) that the shape of the pulse-train solution (25) is similar to that for the pulse-train solution (34) for the value of the elliptic modulus $k=0.999$.
-The key observation is that the amplitudes, widths and wavenumbers of the derived pulse-train solutions are different whereas their velocity is the same.
-More importantly, the relation (6) shows that the velocity of the pulse-trains can be significantly reduced for suitable parameters of the waveguiding media, which may find application in developing slow light systems.
-In Fig.
-5, we present the dependence of velocity on the group velocity dispersion parameter $\alpha$ of the fiber for three different values of the quartic dispersion parameter $\epsilon$.
-It can clearly be seen that the velocity decreases intensely in the magnitude with the increase of dispersion parameter $\alpha$.
+5, we present the dependence of velocity on the group velocity dispersion parameter $\alpha$ of the medium for three distinct values of quartic dispersion coefficient $\epsilon$.
+We clearly see that the wave velocity decreases intensely in the magnitude as the value of dispersion parameter $\alpha$ is increased.
 The decrease of velocity implies that the slowing motion of pulse-trains is possible in the transmitting medium when considering large dispersion parameter $\alpha$.
 It is also observed that the pulse velocity has large magnitude for small negative values of $\epsilon$.
-Hence, one concludes that the velocity of pulse-trains can be controlled by judicious choice of the dispersion parameters.
+Hence, one concludes that the velocity of pulse-trains can be controlled by judicious choice of the dispersion coefficients.
 III.2 Quartic and dipole soliton solutions
@@
 In such a case, the periodic solution (16) degenerates to the quartic soliton solution [35]:
-$\psi(z,\tau)=P_{0}~{}\mathrm{sech}^{2}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[i(\kappa z-\delta\tau+\theta)],$ (36)
+$\psi(z,\tau)=P_{0}~\mathrm{sech}^{2}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (36)
 where the inverse temporal width $w_{0}$, amplitude $P_{0}$ and wave number $\kappa$ of the soliton pulse are given by
@@
 $P_{0}=\pm\sqrt{\frac{-3}{10\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (38)
-$\kappa=-\frac{4}{25\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (39)
-with $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$. Also the velocity $v$ and frequency shift $\delta$ in this quartic soliton solution are defined by Eqs. (5) and (6).
-In addion, when $k\rightarrow 1$, the periodic solution (25) transforms into our dipole soliton solution [7]:
-$\psi(z,\tau)=P_{0}~{}\mathrm{sech}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\mathrm{th}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[i(\kappa z-\delta\tau+\theta)],$ (40)
-where the inverse temporal width $w_{0}$, amplitude $P_{0}$ and wave number $\kappa$ of the soliton pulse take the form
+$\kappa=-\frac{4}{25\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right),$ (39)
+with $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$. Here the soliton frequency shift $\delta$ is given by Eq. (5) and the velocity $v$ is defined by Eq. (6).
+Notice that for this quartic soliton waveform, the relation between the amplitude and duration is given by
+$P_{0}\tau_{0}^{2}=-2\sqrt{\frac{-30\epsilon}{\gamma}}.$ (40)
+provided that $\gamma\epsilon<0$.
+In addition, when $k\rightarrow 1$, the periodic solution (25) transforms into our dipole soliton solution [36]:
+$\psi(z,\tau)=P_{0}~\mathrm{sech}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\mathrm{th}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (41)
+where the inverse temporal width $w_{0}$, amplitude $P_{0}$ and wave number $\kappa$ of the soliton pulse take the form,
 $w_{0}=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}}},$ (42)
 $P_{0}=\pm\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (43)
-$\kappa=\frac{11}{100\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (43)
-with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$. The velocity $v$ and frequency shift $\delta$ in this dipole soliton solution are given by Eqs. (5) and (6).
-It is interesting to note that if one considers the limiting case with $k=1$, the periodic wave solution (34) also degenerates to the same soliton-like solutions (40).
-IV Conclusion
+$\kappa=\frac{11}{100\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right),$ (44)
+with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$. Notice that the frequency shift $\delta$ of this dipole soliton is given by Eq. (5) and its velocity $v$ is defined by Eq. (6).
+It is interesting to note that if one considers the limiting case with $k=1$, the pulse-train solution (34) also becomes the same dipole-type soliton solution (41).
+Notice that for this kind of solitons, we obtain a relation between the pulse amplitude and duration as:
+$P_{0}\tau_{0}^{2}=2\sqrt{\frac{30\epsilon}{\gamma}}.$ (45)
+provided that $\gamma\epsilon>0$.
+IV Propagation of pulse-trains in extended NLSE with higher-order dispersion
+For completeness, we have numerically simulated the propagation of pulse-train solutions (16), (25) and (34) in Eq.
+(1) by utilizing the split-step Fourier method [52].
+The physical parameters used in this model for the case of pulse-train solution (16) are representative of a femtosecond Ti: sapphire laser [53]: $\beta_{2}=-100$ fs2/cm, $\beta_{3}=-500$ fs3/cm, $\beta_{4}=-5000$ fs4/cm, and $\gamma=2.36\times 10^{-6}$ W-1cm-1.
+Note that these experimental parameters satisfy the relations $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$ which are necessary and sufficient for the existence of the pulse-train solution (16).
+As concerns the pulse-train solutions (25) and (34), we use the parameters of the silicon-based slot waveguide as [36, 41]: $\beta_{2}=-0.05$ ps2/m, $\beta_{3}=1.9\times 10^{-5}$ ps3/m, $\beta_{4}=2.5\times 10^{-5}$ ps4/m, and $\gamma=42$ W-1cm-1.
+We point out that with these system parameters, the relations $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$ for the existence of pulse-train solutions (25) and (34) are fulfilled.
+Figure 6 illustrates the evolutions of periodic pulse-trains propagating through a distance of $20$ dispersion lengths.
+Here the elliptic modulus is chosen as $k=0.9$.
+We can see that the pulse-train solutions (16), (25) and (34) can preserve their shape during propagation due to the balance between the effects of quadratic, cubic and quartic dispersions and the Kerr nonlinearity in the system.
+Finally, the figures presented above indicate that the newly found pulse-train solutions have no oscillated tails on the wings of pulses.
+It important to mention that experimental and numerical studies have demonstrated that solitons propagating in nonlinear Kerr media with dominant fourth-order dispersion are approximately Gaussian with oscillating tails [54, 55].
+By considering the influence of both second-, third- and fourth-order dispersions on the pulse propagation through fibers, one of the present authors has recently obtained quartic solitons with a $\mathrm{sech}^{2}$ shape which does not carry oscillating tails [35].
+Under these conditions, we have also found that dipole solitons without oscillating tails are formed in the highly dispersive fiber system [36].
+These important results show that dispersive effects up to the fourth order and Kerr nonlinearity can achieve a perfect balance under certain conditions and lead to the formation of solitons without oscillating tails.
+No doubt, the pulse-trains presented here may be of guidance for the experimental generation of optical periodic waves due to their precise nature and interesting amplitude-duration relationship.
+V Stability of the pulse-train solutions
+[…]
+Figure 7: The numerical evolution; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34) under the perturbation of white noise with maximal value as $0.1$.
+The values of parameters are the same as in Figs.
+2 (a), (b) and (c) respectively.
+VI Conclusion
 In conclusion, we have discovered three classes of pulse-trains whose amplitudes, widths and wavenumbers are different with the distinctive property of exhibiting a relation between the amplitude and duration determined by the sign of a joint parameter uniquely.
@@
 It is also shown that the pulse-trains are obtained without necessarily assuming a specified condition on the fiber parameters.
-The results presented here may be useful for experimental realization of pulse-trains in optical fibers and may help in stimulating more research in understanding their optical transmission properties.
+Due to the precise nature, the nonlinear waves presented here may be profitably used in designing the optimal fiber system experiments.
+These results can be exploited to experimentally realize pulse-trains in fiber systems and may help in stimulating more research in understanding their optical transmission properties.